OpenStudy (anonymous):

Find the general solution of x'=(3, 2, -2, -2)x. (this is matrix, 3 and 2 on the left, -2 and -2 on the right.)

OpenStudy (loser66):

\[\left[\begin{matrix}3&2\\-2&-2\end{matrix}\right]\] step 1: find eigenvalue: \[\left[\begin{matrix}3-\lambda&2\\-2&-2-\lambda\end{matrix}\right]\]characteristic equation: \[(3-\lambda)(-2-\lambda)+4=0\]

OpenStudy (anonymous):

How do you find the eigenvalue?

OpenStudy (loser66):

hey, this information comes from linear algebra course. If you didn't take that course, how can you be in differential equation class?

OpenStudy (anonymous):

I suck at Linear Algebra. Sorry.

OpenStudy (loser66):

I didn't finish the first step yet.

OpenStudy (anonymous):

Please continue.

OpenStudy (loser66):

so, do you know how to construct differential equation?

OpenStudy (loser66):

ok, from the equation above, I have the simplified form is \[\lambda^2-\lambda-2=0\\(\lambda -2)(\lambda+1)=0\\\lambda_1=2~~~and~~~ \lambda_2=-1 \] those are eigenvalues of the matrix. From those, plug into the second matrix, which I construct above to find eigenvectors

OpenStudy (loser66):

For \(\lambda_1=-1\) the matrix above (mine) becomes \[\left[\begin{matrix}4&2\\-2 &-1\end{matrix}\right]\], so, just calculate it by \[\left[\begin{matrix}4&2\\-2 &-1\end{matrix}\right] \left(\begin{matrix}a\\b\end{matrix}\right)=\left(\begin{matrix}0\\0\end{matrix}\right)\]

OpenStudy (loser66):

so, -2a =b, if I chose a =1, so b = -2. therefore \[\left(\begin{matrix}1\\-2\end{matrix}\right)\]is eigenvector corresponds to eigenvalue -1 Can you imitate my stuff to find out the eigenvector corresponds to eigenvalue 2? We are near final answer, girl. PRACTICE, Please

OpenStudy (anonymous):

How did you get -2a=b?

OpenStudy (loser66):

from the matrix |4 2 | | -2 -1| you can pick the last row time with vector (a,b) that mean -2a -b=0 or -2a =b

OpenStudy (loser66):